The terms of the sale were 4/3, net 76. What is the effective annual rate of interest?

Convert your answer to percentage and round off to two decimal points. Do not enter % in the answer box.
Answer #1

The interest rate for the term of the discount is: Interest rate = 0.04/0.96 = 4.17% Discount period = 76-3 = 73 days EAR = E